# Slot-BERT: Self-supervised Object Discovery in Surgical Video

### v2

@@ 4.8 Adaptation to longer sequences
| Short (5 frames) | Long (11 frames) | Endovis | Thoracic | ||||||||
@@ 4.10 Experiment with slot-Mixer Decoder
To evaluate our method with the alternative Slot-Mixer decoder (Zadaianchuk et al., 2023), we conducted experiments using the Slot-BERT framework and compared it against two additional configurations. The results are reported for the MICCAI benchmark, where the model is trained and tested on MICCAI, and the zero-shot Endovis benchmark, where the model trained on MICCAI is tested directly on Endovis without fine-tuning. The first configuration, Mixer Only, combines a vanilla RNN Slot Attention with the Slot-Mixer for feature reconstruction. The second, Contrast+Mixer, incorporates slot contrastive loss into the training process while still relying on the Slot-Mixer for decoding. Finally, Mixer-full represents the full proposed method.
@@
These findings confirm that our method is able to surpass the vanilla RNN-based baseline using either MLP broadcast or Slot-Mixer decoder, achieving high localization and segmentation performance. For more qualitative results refer to the section F.1 of supplementary materials.
@@ A Illustration of using TST module for next slot initialization
B Experiment with different slot contrastive temperatures
We conducted experiments to evaluate the impact of the slot contrastive temperature () in the calculation of contrastive loss on segmentation performance in both same-domain and zero-shot transfer scenarios. The model was trained on the MICCAI dataset and tested on both the MICCAI test set (same-domain performance) and the unseen EndoVis dataset (zero-shot transfer performance).
The results, presented in Table S1, demonstrate the nuanced role of contrastive temperature in influencing segmentation accuracy. In the same-domain context, a temperature setting of achieved the best mBO-V score of 49.6%, along with a frame-level accuracy (mBO-F) of 52.8%. Meanwhile, resulted in the highest mBO-F score on the MICCAI data (53.0%). This suggests that a smaller temperature effectively balances inter-slot distinctiveness while maintaining smooth transitions across highly similar instrument categories, which is a common characteristic of the MICCAI dataset, where some instruments share significant visual similarities.
In the zero-shot transfer setting tested on the EndoVis dataset, the performance trends shifted. Here, achieved the highest mBO-V score of 44.0% and an mBO-F of 48.4%. This indicates that a slightly higher temperature helps the model better separate and distinguish slots in a dataset with more varied instrument appearances, improving generalization. Interestingly, while yielded comparable performance in the zero-shot scenario (mBO-V: 43.6%, mBO-F: 47.5%), it decreased the segmentation accuracy compared to , suggesting that overly increasing the temperature can negatively impact performance.
These findings emphasize the importance of carefully selecting the contrastive temperature based on the target domain and task requirements. The contrastive temperature provides flexibility, enabling the model to adjust its behavior for optimal performance in both same-domain and zero-shot settings.
| Same domain | Zeroshot transfer | ||||
| mBO-V (%) | mBO-F (%) | mBO-V (%) | mBO-F (%) | ||
| 1.2 | 48.5 ± 0.5 | 52.2 ± 0.3 | 43.6 ± 0.4 | 47.5 ± 0.3 | |
| 1.1 | 48.9 ± 0.2 | 52.8 ± 0.2 | 44.0 ± 0.2 | 48.4 ± 0.1 | |
| 1.0 | 48.2 ± 0.2 | 51.9 ± 0.3 | 43.5 ± 0.3 | 47.6 ± 0.3 | |
| 0.8 | 49.6 ± 0.1 | 52.8 ± 0.3 | 43.4 ± 0.3 | 47.3 ± 0.2 | |
| 0.5 | 48.4 ± 0.3 | 53.0 ± 0.2 | 43.0 ± 0.2 | 47.7 ± 0.1 | |
| 0.1 | 48.9 ± 0.5 | 52.6 ± 0.7 | 44.0 ± 0.2 | 47.5 ± 0.2 | |
@@ E Additional results on transfer learning on Endovis and Thoracic data
F Additional qualitative results
F.1 Comparison on using MLP and slot-Mixer decoder
@@
Both MLP and Slot-Mixer decoders demonstrate unique strengths. With the MLP decoder, the segmentation masks tend to cover objects more comprehensively and accurately locate them. On the other hand, the Slot-Mixer decoder exhibits superior temporal consistency. For instance, in the first and third rows of Figure S3, the instrument undergoes significant motion (e.g., moving across the image within a few frames). While the MLP decoder loses track of the object during such maneuvers, the Slot-Mixer decoder successfully maintains object tracking. This is evident in the image sequence, where the Slot-Mixer preserves the correct mask order despite the large motion, showcasing its robustness in maintaining temporal coherence.
F.2 Zero-shot Segmentation of longer video sequence

### v3

@@ 4.8 Adaptation to longer sequences
| Method | mBO-V (%) | mBO-F (%) | FG-ARI (%) | Corloc (%) |
| DINO-Saur(Seitzer et al., 2023) | 30.96 ± 0.29 | 42.15 ± 0.09 | 45.11 ± 0.41 | 52.79 ± 0.71 |
| SAVi(Kipf et al., 2022) | 25.39 ± 0.25 | 35.23 ± 0.02 | 35.98 ± 0.01 | 42.56 ± 0.23 |
| STEVE(Singh et al., 2022) | 25.37 ± 0.06 | 33.99 ± 0.01 | 36.83 ± 0.01 | 45.76 ± 0.06 |
| Slot-Diffusion(Wu et al., 2023b) | 30.67 ± 0.38 | 40.57 ± 0.15 | 43.96 ± 0.16 | 52.94 ± 0.22 |
| Video-Saur(Zadaianchuk et al., 2023) | 42.31 ± 0.09 | 53.04 ± 0.07 | 56.81 ± 0.32 | 68.68 ± 0.39 |
| Slot-BERT (ours) | 44.37 ± 0.27 | 55.31 ± 0.09 | 61.68 ± 0.25 | 73.22 ± 0.24 |
@@
To further demonstrate performance on even longer temporal context windows, we train and test Slot-BERT on 30 second episodes. The results and comparison to other methods are shown in Table 6. Note that, compared to shorter videos (e.g., 11 frames), the video-level mBO score decreased by around 2%. Nevertheless, Slot-BERT beats state-of-the-art approaches across all metrics and maintains a high image-level segmentation accuracy in comparison to training on shorter clips. Note that training on longer sequences is not necessarily beneficial once the temporal structure becomes less informative about frame-to-frame transitions. An optimal training protocol for Slot-BERT would thus take into account the length and complexity of typical surgical actions and tasks.
@@
| 5 frames | 11 frames | Endovis | Thoracic | ||||||||
@@ 4.9 Ablation study
To illustrate the effects of the TST module and the contrastive loss, we compute the cosine similarity between slot representation vectors and visualize them as heat maps across frame sequences (Figure 5). When the TST module is disabled, temporal consistency deteriorates. As shown on the left side of the figure, the mask originally tracking the instrument on the left (slot index 0, blue mask) switches to another slot (index 2, green mask). Similarly, the tissue at the bottom, initially represented by a single slot (index 1), is later fragmented into two slots (indices 1 and 3). These inconsistencies are also evident in the similarity matrix, where pronounced fluctuations appear in the rows corresponding to these slots. In contrast, under the full Slot-BERT configuration, object tracking is more stable. For example, the same left instrument (slot index 6, purple mask) remains consistently represented by a single slot across the entire sequence, which is reflected by the smoother, more stable row in the similarity matrix at the bottom.
The effect of removing the contrastive loss is shown on the right of Figure 5. In this setting, slot similarities remain uniformly high (above 0.6), whereas with contrastive loss they are pushed lower, clustering around zero, enforcing that slots encode diverging directions in the representation hyper-space. Without contrastive loss, the model exhibits segmentation instability: two distinct instruments in the lower-right region of the frame are merged into a single slot (index 4) across the video. This ambiguity is reflected in the similarity matrix as bright regions indicating high correlation between slot vectors. By contrast, when contrastive loss is applied, slots are explicitly encouraged to diverge, enabling the two neighboring instruments to be represented by separate slots (indices 1 and 3, shown in orange and red masks).
4.10 Contrastive hyperparameter tuning analysis
To investigate the effect of contrastive loss hyper-parameters on model performance, we systematically varied the contrastive loss weight () from 0.0001 to 0.1 and the temperature parameter () from 0.05 to 100. Figure 6 visualizes the tendency.
Overall, we observe that a larger contrastive weight (e.g., ) requires a higher temperature to achieve better accuracy. This is likely because over-enforcing a strong dissimilarity between slots can have negative impact on object-centric learning; a higher temperature effectively smooths the contrastive distribution, mitigating the over-penalization caused by large weights. Conversely, when the contrastive weight is too small (e.g., ), the effect of varying the temperature is marginal, as the relative change of influence on training dynamics is much less for this combination.
These findings suggest that a moderate weight on the contrastive objective, coupled with a balanced temperature scaling, provides the best trade-off between encouraging slot diversity and maintaining stable optimization. In particular, tuning the parameters to and yields the highest video-level segmentation performance, achieving 50.1 mBO-V, which represents a notable improvement over the baseline model without contrastive learning (46.3 mBO-V).
@@ 4.11 Experiment with slot-Mixer Decoder
To evaluate our method with the alternative Slot-Mixer decoder (Zadaianchuk et al., 2023), we conducted experiments using the Slot-BERT framework and compared it against two additional configurations. The results are reported for the MICCAI benchmark, where the model is trained and tested on MICCAI, and the zero-shot Endovis benchmark, where the model trained on MICCAI is tested directly on Endovis without fine-tuning. The first configuration, Mixer Only, combines a vanilla RNN Slot Attention with the Slot Mixer for feature reconstruction. The second, Contrast+Mixer, incorporates slot contrastive loss into the training process while still relying on the Slot-Mixer for decoding. Finally, Mixer-full represents the full proposed method.
@@
These findings confirm that our method is able to surpass the vanilla RNN-based baseline using either MLP broadcast, or Slot-Mixer decoder which improves decoding quality by introducing interactions between slots, achieving high localization and segmentation performance. For more qualitative results refer to the section G.1 of supplementary materials.
@@ 4.13 Experiment with different slot numbers
4.14 Computational efficiency analysis
We benchmark the computational efficiency of Slot-BERT against existing object-centric video models. The forward inference time per frame is measured on a single NVIDIA RTX A6000 GPU using video clips of seconds, where we set the temporal context window to seconds. We also report the floating point operations per second (GFLOPs), which quantify the computational cost of processing a single frame. The slot number for all methods is set to 7 in this experiment.
As shown in Table 9, Video-Saur achieves the fastest inference speed ( ms per frame), while Slot-BERT maintains a competitive runtime efficiency ( ms per frame) with only a minor computational overhead compared to other baselines. Although Slot-Diffusion reports the lowest GFLOPs (), its inference speed is not competitive due to the iterative nature of diffusion-based decoding process.
We also evaluated the memory consumption under real-time video processing with slot mixer decoder. In this setting, the RNN-based model, Video-Saur requires MB. By incorporating TST module With a 5-frame context window, the overhead of slot-BERT is only MB ( MB total), and with a 30-frame context window, the overhead remains modest at MB, thanks to the latent space operation of temporal attention.
@@ 5.2 Future Directions
Beyond incorporating additional modalities such as optical flow or depth maps to refine boundary accuracy, another promising direction is domain-specific high-resolution fine-tuning. This could be achieved by leveraging models pretrained on lower-resolution inputs (as typically required by default backbones like DINO) as anchors for adaptation (Didolkar et al., 2025). In addition, DINOv3 (Siméoni et al., 2025) that is already trained with a high-resolution adaptation phase supports effective inference using high-resolution features.
@@ A Illustration of using TST module for next slot initialization
B Visualization of channel-wise object-centric attention overlay
We provide extended visualizations of the slot attention maps across independent channels (Figure S2). Following the presentation style of Figure 5, to indicate correspondence, the slot index number is overlaid in white on each colored mask of the first frame. In this example, Slot-BERT is trained with seven slots for surgical videos, where these slots are allocated to attend to different instruments and tissue regions.
@@ E Additional results on transfer learning on Endovis and Thoracic data
F Additional Results on Non-Surgical Data
We additionally evaluate Slot-BERT on non-surgical datasets, including the natural real-world dataset YT-VIS (Xu et al., 2019) and the synthetic dataset MOVi-E (Greff et al., 2022). Following prior work (Wu et al., 2023b; Zadaianchuk et al., 2023), we use slots for training on YT-VIS and slots for MOVi-E. For evaluation, we sample frames on YT-VIS and frames on MOVi-E, consistent with the setup in Video-Saur (Zadaianchuk et al., 2023).
As shown in Table S3, on the real-world YT-VIS dataset our method outperforms the second-best Video-Saur by over mBO-V. On the synthetic MOVi-E dataset, Slot-BERT achieves comparable FG-ARI performance to Video-Saur. We note that MOVi-E provides less challenging conditions for object discovery due to its fixed camera viewpoint and relatively simple backgrounds which likely reduces the performance gap between methods.
G Additional qualitative results
G.1 Comparison on using MLP and slot-Mixer decoder
@@
Both MLP and Slot-Mixer decoders demonstrate unique strengths. With the MLP decoder, the segmentation masks tend to cover objects more comprehensively and accurately locate them. On the other hand, the Slot-Mixer decoder exhibits superior temporal consistency , since its within-frame attention at decoding provides synergistic benefits when combined with the temporal attention (TST), further stabilizing segmentation across video frames. For instance, in the first and third rows of Figure S4, the instrument undergoes significant motion (e.g., moving across the image within a few frames). While the MLP decoder loses track of the object during such maneuvers, the Slot-Mixer decoder successfully maintains object tracking. This is evident in the image sequence, where the Slot-Mixer preserves the correct mask order despite the large motion, showcasing its robustness in maintaining temporal coherence.
G.2 Zero-shot Segmentation of longer video sequence